-(x-2)+4x=5x-8-2(x-5)

Answer by stanbon(75887)

-(x-2)+4x=5x-8-2(x-5)
------------
-x + 2 + 4x = 5x -8 -2x + 10
-------
3x+2 = 3x+2
---
True for all Real Numbers
=================================
Cheers,
Stan H.